Two objects A and B having inertias Ma=m and Mb=1.1m are moving with velocities Va=V and Vb=8.4v.
Find the velocities of object B in the zero momentum frame of reference
1
Expert's answer
2020-05-19T10:52:09-0400
"v_a'=v_a-\frac{m_av_a+m_bv_b}{m_a+m_b}"

"v_a'=v-\frac{mv+1.1(8.4)mv}{m+1.1m}=-3.9v"

"v_b'=8.4v-\frac{mv+1.1(8.4)mv}{m+1.1m}=3.5v"